Explain income elasticity of demand and its measurement in case of inferior goods and superior goods with table and diagrams


1
Expert's answer
2021-02-02T09:28:12-0500

Let's start with definitions. The income elasticity of demand measures the relationship between a change in quantity demanded for given good and a change in real income. The income elasticity can be calculated as follows:


"E_y=\dfrac{\%\ Change\ in\ demand}{\%\ Change in\ income}."

An inferior goods are the type of goods whose demand decreases when the income rises. Therefore, the income elasticity of demand always will be negative. An example of such a good is the demand of cigarettes. The superior goods (or the luxury goods) are the type of goods that displace the demand of inferiors goods after a rise in consumer's income. Therefore, the income elasticity of demand > +1. An example of such a good is the demand of smoked salmon.

Let's describe a case of income elasticity of demand in case of inferior goods. Let's imagine that we have the following table for good X:



According to our formula, we can calculate the income elasticity of demand for good X:


"E_y=\dfrac{\dfrac{Q_2-Q_1}{Q_1}}{\dfrac{Y_2-Y_1}{Y_1}}=\dfrac{\dfrac{80-200}{200}}{\dfrac{80000-40000}{40000}}=-0.6"


Let's describe a case of income elasticity of demand in case of superior goods. Let's imagine that we have the following table for good Y:


According to our formula, we can calculate the income elasticity of demand for good Y:


"E_y=\dfrac{\dfrac{Q_2-Q_1}{Q_1}}{\dfrac{Y_2-Y_1}{Y_1}}=\dfrac{\dfrac{165-150}{150}}{\dfrac{108000-100000}{100000}}=+1.25"